# The Jaynes-Cummings model as an autonomous Maxwell demon

### v2

@@ Abstract
Autonomous feedback and qubit purification. Remarkably, the behavior of the machine presents a third regime on an even longer timescale $t\sim\sqrt{n_{0}}g^{-1}$, in which the cavity performs an effective drive depending on the result of the measurement in the {|±$\{|\pm_{y}\rangle\}$ basis, thereby performing an autonomous feedback. This feedback is able to purify any initial qubit state when $n_{0}$ is sufficiently large [40]. We first note that at $t\sim t_{c}$, the cavity field $\text{Tr}\{a\hat{\rho}_{C}^{\nu\nu}\}=\alpha_{\nu}(t)$ depends on the measurement result $\nu$. It is instructive to analyze the joint qubit-cavity dynamics along the two independent branches associated with the qubit found in |±$|\pm_{y}\rangle$. As, inside each branch, the qubit-cavity state is factorized up to ${\cal O}(1/n_{0})$ corrections, one can again consider the cavity as an ideal work source, inducing a classical drive on the qubit of the form $\hat{H}^{\text{eff}(\nu)}_{Q}(t)=i\hbar g(\alpha_{\nu}(t)\sigma_{+}-\alpha_{\nu}^{*}(t)\sigma_{-})$ [32], where, unlike early times ($t<t_{c}$), the effective driving Hamiltonian depends on the measurement result. At $t_{c}$, the two Hamiltonians almost coincide with $\hat{H}_{Q}^{\text{eff}}(0)$, and the qubit measured state is aligned with the field, whatever $\nu$. Then, two different field axes emerge. From Eq. (4), we see that they both lie in the equator and rotate in opposite directions with angular velocities ±$\pm\frac{g}{2\sqrt{n_{0}}}$ much smaller, by a factor of $1/n_{0}$, than the energy splitting of $\hat{H}^{\text{eff}(\nu)}_{Q}(t)$, still given by the Rabi frequency $\Omega$. As a consequence, the qubit follows adiabatically the rotation axis inside each measurement branch (see Fig. 3). The two field axes meet again at $t=t_{\text{min}}=\sqrt{n_{0}}\pi g^{-1}$ where they overlap with the $x$-axis of the Bloch sphere, driving in both branches the qubit to state $|+_{x}\rangle$. As a consequence of this autonomous feedback, an arbitrary initial qubit state $\hat{\rho}_{Q}(0)$ is deterministically purified and brought to $|+_{x}\rangle$ (or generally for a phase $\phi_{0}$ to state $(1/\sqrt{2})\left(|g\rangle+e^{i\phi_{0}}|e\rangle\right)$) at time $t_{\text{min}}$.
@@
As seen in Fig. 1(a), the entropy of the qubit goes through a series of minima with a decreasing purity at times $t_{k}\simeq(2k+1)\sqrt{n_{0}}\pi$ (with $t_{1}\equiv t_{\text{min}}$), except when it starts already in one of the two pointer states |±$|\pm_{y}\rangle$ (red curve). In this case, the cavity-qubit correlations associated with the measurement do not build up, and the qubit entropy increase is only due to intra-branch correlations associated with finite $n_{0}$ behavior. Between two consecutive minima, the qubit goes through entropy maxima of increasing value. After a few oscillations, as intra-branch qubit-cavity correlations keep growing, the qubit entropy saturates at $\log 2$, the maximum physical value for a qubit.
@@
Thermodynamic signature of an autonomous Maxwell demon – We now use our thermodynamic framework to analyze the purification process, and show that it can be interpreted as an autonomous Maxwell demon. We assume that the qubit is initially prepared in the maximally mixed state $\rho_{Q}(0)=\mathds{1}/2$. Fig. 4 shows the evolution of the qubit and cavity entropies, together with their quantum mutual information $I_{CQ}(t)=S_{C}(t)+S_{Q}(t)-S_{CQ}(t)$. We also plotted the work $W_{C}(t)$ and heat $Q_{C}(t)$ provided by the cavity, together with its internal energy variation $\Delta E_{C}$. We first note that the machine is able to decrease the qubit entropy (initially equal to its maximal value $\log 2$) down to a minimum value reached at time $t_{\text{min}}$, at the expense of work performed by the cavity, akin to a Szilard engine [42, 43].
@@
Figure 4: (a) von Neumann entropy of the cavity $S_{C}$ and of the qubit $S_{Q}$, and mutual information between the qubit and cavity $I_{CQ}$. (b) Cavity heat $Q_{C}$, work $W_{C}(t)$ and internal energy variation $\Delta E_{C}(t)$. For both (a) and (b), $n_{0}=50$, $\bar{n}=1$. Plotted quantities are computed from exact calculation with truncated cavity Hilbert space.
We then build on the dynamical analysis above to interpret the evolution of the thermodynamic quantities associated with the purification process. As the qubit-cavity correlations associated with the measurement grow during time interval $[0,t_{c}]$, the mutual information increases from 00 (as expected for an initial factorized state) to almost its maximal value $\log 2$ reached around $t_{c}$. The reduced entropy of the cavity also increases by the same amount. This step corresponds to the transfer of information about the qubit into the demon’s memory. Over the interval $[t_{c},t_{\text{min}}]$, the autonomous feedback of the cavity induces conditional rotations of the qubit state in the equator of the Bloch sphere based on the qubit state. This process is able to decrease the qubit entropy, but has no work cost (as seen on Fig. 4(b)) as the qubit’s internal energy remains constant. This apparent violation of the second law of thermodynamics is enabled by the consumption of the mutual information built during the measurement stroke, as can be seen in Fig. 4(a). Consumption of mutual information for an effect analogous to work expenditure is precisely the signature of a Maxwell demon [44, 43, 45]. This behavior here emerges autonomously from the Jaynes-Cummings interaction, and is naturally captured by our thermodynamic framework. We finish by noting that the evolution over the time intervals $[t_{k},t_{k+1}]$, correspond to a new measurement step followed by another feedback step, with degraded performances due to the building up of intra-branch correlations.
@@
In conclusion, we have analyzed the dynamics of a qubit coupled to a cavity prepared in a displaced coherent state of very large displacement amplitude, a state commonly considered as guaranteeing a classical drive behavior of the cavity. Owing to our framework for the thermodynamics of autonomous machines, we have shown that at short enough times, the cavity indeed behaves as an ideal work source inducing a time-dependent drive on the qubit. However, over a time-scale set by the qubit-cavity interaction strength, the cavity performs a measurement of the qubit state in a basis set by the initial cavity displacement phase, responsible for the collapse of the Rabi oscillations. We have shown that this measurement is followed by an autonomous feedback, a state-dependent drive able to purify the qubit. From the analysis of thermodynamic quantities, we relate this purification to the cavity behaving as an autonomous Maxwell demon, spending work to build mutual information with the qubit, and then consuming this mutual information decrease the qubit’s entropy. Those results demonstrate the potential of consistent thermodynamic frameworks, like the one of Ref. [30], to shed new light on complex quantum dynamics. The parameters needed to implement our scheme are within state-of-the-art, e.g. in superconducting qubit-cavity systems, where strong couplings with respect to damping rates can be achieved. Natural extensions of this work include optimizing the initial cavity state to improve the purification process, e.g. by attaining purer states at a fixed feedback time, and testing the robustness of the purified qubit state to cavity preparation and residual qubit-cavity detunings.
@@
Acknowledgements – C.E. acknowledges funding from French National Research Agency (ANR) under project ANR-22-CPJ1-0029-01. We thank Alexia Auffèves for fruitful discussions.
Supplemental Material: The Jaynes-Cummings model as an autonomous Maxwell demon
I A. Physical Model
@@
$\hat{H}$ $=\hat{H}_{C}+\hat{H}_{Q}+\hat{V}_{QC}=\hbar\omega_{0}a^{\dagger}a+\hbar\omega_{0}\sigma_{+}\sigma_{-}+i\hbar g\left(a\sigma_{+}-a^{\dagger}\sigma_{-}\right)$ (S1)
@@
Due to the resonance condition, the evolution of the total system in the interaction picture is governed by Hamiltonian $\hat{H}_{I}(t)=\hat{V}_{\text{QC}}$. We deduce the interaction picture evolution operator $\hat{U}(t)=e^{-\frac{it}{\hbar}\hat{V}_{QC}}$, which can expanded as [35]:
$\hat{U}\left(t\right)=\cos\left(gt\sqrt{a^{\dagger}a+1}\right)|e\rangle\langle e|+\cos\left(gt\sqrt{a^{\dagger}a}\right)|g\rangle\langle g|+\frac{\sin\left({gt\sqrt{a^{\dagger}a+1}}\right)}{\sqrt{a^{\dagger}a+1}}a\sigma_{+}-\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}\sigma_{-}$ (S2)
II B. Quasi-ideal work source behavior of the cavity
@@
$\hat{\rho}\left(0\right)=|e\rangle\langle e|\otimes D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right),\hskip 5.0pt\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}\;.$ (S3)
@@
$\hat{\rho}\left(t\right)$ $=|e\rangle\langle e|\otimes\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)$ (S4)
-- $|e\rangle\langle g|\otimes\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}$
$-|g\rangle\langle e|\otimes\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)$
$+|g\rangle\langle g|\otimes\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}\;.$
@@
$\hat{\rho}_{Q}\left(t\right)=\begin{bmatrix}P_{e}\left(t\right)&C_{eg}\left(t\right)\\ C_{eg}\left(t\right)&1-P_{e}\left(t\right)\end{bmatrix}$ (S5)
@@
$P_{e}\left(t\right)$ $=\sum_{n}\cos^{2}\left({gt\sqrt{n+1}}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ (S6)
$C_{eg}\left(t\right)$ $=-\sum_{n}\cos\left({gt\sqrt{n+2}}\right)\sin\left({gt\sqrt{n+1}}\right)\langle n+1|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle\;.$
For $n_{0}\gg 1$ the terms $\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ and $\langle n+1|D\left(\alpha\right)W_{\beta\left(0\right)}D^{\dagger}\left(\alpha\right)|n\rangle$ as a function of $n$ are close to Gaussian distributions peaked around $n_{0}$ with variance much smaller than the mean. This allows us to expand the trigonometric functions in $P_{e}\left(t\right)$ and $C_{eg}\left(t\right)$ around $n_{0}$. In the following, we parameterize the population and coherence of qubit in terms of rotation angle $\theta=2gt\sqrt{n_{0}}$ in the Bloch sphere.
@@
$P_{e}\left(\theta\right)$ $=\sum_{n}\cos^{2}\left({\frac{\theta}{2}\sqrt{1+\frac{n-n_{0}+1}{n_{0}}}}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ (S7)
@@
×$\times\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$
@@
$\delta P_{e}\left(\theta,\bar{n}\right)=-\frac{\theta}{16}\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)\;.$ (S8)
@@
$C_{eg}\left(\theta\right)$ $=-e^{i\phi_{0}}[\frac{1}{2}\sin{\theta}+\frac{-2\theta+\left(3+2\bar{n}\right)\theta\cos\theta-\left(1+2\bar{n}\right)\left(1+\theta^{2}\right)\sin\theta}{16n_{0}}+$ (S9)
@@
$=e^{i\phi_{0}}[\frac{1}{2}\sin{\theta}+\frac{\delta C_{eg}\left(\theta,\bar{n}\right)}{n_{0}}+O\left(\frac{1}{n_{0}^{2}}\right)]$
@@
$\delta C_{eg}\left(\theta,\bar{n}\right)=\frac{-2\theta+\left(3+2\bar{n}\right)\theta\cos\theta-\left(1+2\bar{n}\right)\left(1+\theta^{2}\right)\sin\theta}{n_{0}}$ (S10)

### v4

@@ VII Autonomous feedback and qubit purification
Remarkably, the behavior of the machine presents a third regime on an even longer timescale $t\sim\sqrt{n_{0}}g^{-1}$, in which the cavity induces an effective drive on the qubit depending on the result of the measurement in the $\{|\pm_{y}\rangle\}$ basis. This evolution therefore constitutes an example of autonomous feedback. This feedback is able to purify any initial qubit state when $n_{0}$ is sufficiently large, as shown for zero initial cavity temperature in Ref. [43], and here extended to an initially displaced thermal state. We first note that at $t\sim t_{c}$, the cavity field $\text{Tr}\{a\hat{\rho}_{\text{C}}^{\nu\nu}\}=\alpha_{\nu}(t)$ depends on the measurement result $\nu$. It is instructive to analyze the joint qubit-cavity dynamics along the two independent branches associated with the qubit found in $|\pm_{y}\rangle$. As, inside each branch, the qubit-cavity state is factorized up to ${\cal O}(1/n_{0})$ corrections, one can again consider the cavity as an ideal work source, inducing a classical drive on the qubit of the form (see Appendix C.2)
$\hat{H}^{\text{eff}(\nu)}_{\text{Q}}(t)=i\hbar g(\alpha_{\nu}(t)\sigma_{+}-\alpha_{\nu}^{*}(t)\sigma_{-}),$ (24)
where, unlike early times ($t<t_{c}$), the effective driving Hamiltonian depends on the measurement result $\nu$. At $t_{c}$, the two Hamiltonians almost coincide with $\hat{H}_{\text{Q}}^{\text{eff}}(0)$, and the qubit measured state is aligned with the field, whatever $\nu$. Then, two different field axes emerge. From Eq. (21), we see that they both lie in the equator and rotate in opposite directions with angular velocities $\pm\frac{g}{2\sqrt{n_{0}}}$ much smaller, by a factor of $1/n_{0}$, than the energy splitting of $\hat{H}^{\text{eff}(\nu)}_{\text{Q}}(t)$, still given by the Rabi frequency $\Omega$. As a consequence, the qubit follows adiabatically the rotation axis inside each measurement branch (see Fig. 3). The two field axes meet again at $t=t_{\text{min}}=\sqrt{n_{0}}\pi g^{-1}$ where they overlap with the $x$-axis of the Bloch sphere, driving in both branches the qubit to state $|+_{x}\rangle$. As a consequence of this autonomous feedback, an arbitrary initial qubit state $\hat{\rho}_{\text{Q}}(0)$ is deterministically purified and brought to $|+_{x}\rangle$ (or generally for a phase $\phi_{0}$ to state $(1/\sqrt{2})\left(|g\rangle+e^{i\phi_{0}}|e\rangle\right)$) at time $t_{\text{min}}$.
@@
As seen in Fig. 1(a), after $t_{\text{min}}$, the entropy of the qubit goes through a series of minima with a decreasing purity at times
$t_{k}\simeq(2k-1)\sqrt{n_{0}}\pi,\quad(t_{1}\equiv t_{\text{min}}),$ (25)
@@
separated by maxima of increasing entropy. This behavior can be attributed to series of measurement and feedback sequences over the time intervals $t_{k-1},t_{k}]$, with degraded performances due to the accumulation of qubit-cavity correlations during the intra-branch dynamics, at finite $n_{0}$. The entropy maxima are reached after each measurement step. When the qubit starts already in one of the two pointer states $|\pm_{y}\rangle$ (red curve), the measurement has no effect and the cavity-qubit correlations associated with the measurement do not build up, hence the absence of maxima. In that case, the qubit entropy increase is only due to intra-branch correlations. After a few oscillations, as intra-branch correlations keep growing, the qubit entropy saturates at $\log 2$, the maximum physical value for a qubit.
VIII Thermodynamic signature of an autonomous Maxwell demon
@@
We now use our thermodynamic framework to analyze the purification process, and show that the cavity behaves as an autonomous Maxwell demon. We assume in this last section that the qubit is initially prepared in the maximally mixed state $\rho_{\text{Q}}(0)=\mathds{1}/2$ (center of the Bloch sphere). Fig. 4 shows the evolution of the qubit and cavity entropies, together with their quantum mutual information
$I_{\text{QC}}(t)=S_{\text{C}}(t)+S_{\text{Q}}(t)-S_{\text{QC}}(t).$ (26)
We also plotted the work $W_{\text{C}}(t)$ and heat $Q_{\text{C}}(t)$ provided by the cavity, together with its internal energy variation $\Delta E_{\text{C}}$. By comparing the initial and final states of the qubit (respectively, $\rho_{\text{Q}}(0)$ and $|+_{x}\rangle\langle+_{x}|$ in the large $n_{0}$ limit), we first note that the overall effect on the qubit is a net decrease of its entropy from $\log 2$ to 0, without changing its energy.
As it can be seen from the overall evolution of the cavity, this entropy decrease is made possible by an expense of work performed by the cavity on the qubit, as in a Landauer reset protocol [45]. One can use Eq. (7) in the case $j=Q$ and $j^{\prime}=C$, and for the time interval $0,t_{\text{min}}]$ together with the fact that $\Delta E_{C}=-\Delta E_{Q}=0$ (in the large $n_{0}$ limit), to show that $\beta_{C}(0)Q_{C}(t_{\text{min}})=\beta_{C}(0)W_{C}(t_{\text{min}})\geq-\Delta S_{Q}$. The Landauer bound $W_{C}(t_{\text{min}})\geq\beta_{C}^{-1}(0)\log 2$ is therefore retrieved when $n_{0}\gg 1$ such that the purification is complete $\Delta S_{Q}\simeq-\log 2$. We emphasize however that, here, the reset protocol is not occurring via a deterministic driving in presence of a heat source, as typically considered, but rather via the autonomous feedback mechanism induced by the cavity.
@@
Figure 4: (a) von Neumann entropy of the cavity $S_{\text{C}}$ and of the qubit $S_{\text{Q}}$, and mutual information between the qubit and cavity $I_{\text{QC}}$. (b) Cavity heat $Q_{\text{C}}$, work $W_{\text{C}}(t)$ and internal energy variation $\Delta E_{\text{C}}(t)$ associated to the time interval $[0,t]$. For both (a) and (b), $n_{0}=50$, $\bar{n}=1$. Plotted quantities are computed from exact calculation with truncated cavity Hilbert space.
To highlight the signatures of a Maxwell demon, we analyze separately the two time intervals $0,t_{c}]$ (information acquisition) and $t_{c},t_{\text{min}}]$ (purification). We build on the dynamical analysis above to interpret the evolution of the thermodynamic quantities along those two steps. We refer to Fig. 4 where are plotted the numerical evaluations of exact expressions for the relevant information theoretic and thermodynamic quantities. During time interval $0,t_{c}]$ ($gt\in[0,1]$), the qubit-cavity quantum mutual information grows from 0 (as expected for an initial factorized state) to almost its maximal value $I(t_{c})\simeq\log 2$ reached around $t_{c}$ (see Fig. 4a and Appendix C.3 for an analytical derivation in the large $n_{0}$ limit). This step corresponds to the transfer of information about the qubit into the demon’s memory. The reduced entropy of the cavity also increases by the same amount. In the considered case of a qubit starting in the center of the Bloch sphere $\rho_{Q}(0)=\mathds{1}/2$, no Rabi oscillation occur, and the qubit remains in state $\rho_{Q}(0)=\mathds{1}/2$. As a consequence, the qubit’s energy, and therefore the cavity’s energy as well, are constant. The time interval $0,t_{c}]$ is then associated to a work-to-heat transfer inside the cavity, as seen from Fig. 4b.
During time interval $t_{c},t_{\text{min}}]$ ($gt\in[1,gt_{\text{min}}]$), the qubit is purified $S_{Q}(t_{\text{min}})\simeq 0$ (in the large $n_{0}$ limit), while its energy is conserved (its state stays within the equatorial plane of the Bloch sphere on the whole interval, as discussed in Section C.2). Along this time interval as well, the cavity energy is therefore conserved. As the cavity entropy is also constant (see Appendix C.3), the cavity provides no work nor heat to the qubit in the time interval $t_{c},t_{\text{min}}]$. This leads to $\Delta S_{Q}-\beta_{C}(t_{c})Q_{C}=\Delta S_{Q}<0$, which constitutes an apparent violation of Clausius inequality, i.e. the signature of the Maxwell demon. Consistency with the second law is restored by taking into account the resource being consumed, as done in Eq. (8), that is, the quantum mutual information between the cavity and the qubit which simultaneously decreases from $\log 2$ to 0. Those results are analytically confirmed by our perturbative calculation in the large $n_{0}$ limit (see Appendix C.3). We note that the Gaussian approximation used in the unitary regime does not apply, even at dominant order, as the reduced cavity state is a mixture of two orthogonal distributions in phase space, which is highly non-Gaussian. An interesting outcome of the analytical calculation is to show that the inequality (8) is saturated in our model in the large $n_{0}$ limit (see Appendix C.3), meaning a reversible conversion of mutual information into entropy decrease.
As it can be seen in Fig. 4, when $n_{0}$ is finite, the qubit entropy still decreases, albeit to a finite value, and the cavity still exchanges no work nor heat during the interval $t_{c},t_{\text{min}}]$. The entropy decrease of the qubit remains entirely fueled by the consumption of quantum mutual information. We find that such decrease of the qubit’s entropy persists for small values of $n_{0}$, down to a threshold of approximately $n_{0}\sim 5$. Below, the qubit entropy starts exhibiting a non-monotonous evolution and the measurement and feedback steps are not clearly marked anymore.
Thus, we can conclude that a behavior of autonomous Maxwell demon emerges from the Jaynes-Cummings interaction, in the case of an initially coherently-displaced cavity, and is naturally captured by our thermodynamic framework.
We finish by a few comments to discriminate the mechanism of autonomous Maxwell demon presented here from other spontaneous purification processes. Purity minima are expected to occur in any quantum system initialized in a pure state and coupled to another finite-size system. Indeed, Poincaré recurrences [46, 47] ensure that the total system will repeatedly reach states arbitrarily close to its initial state. In contrast, here, the initial state of the qubit is not restored when its purity reaches its minima. Instead, the pure state it reaches is entirely determined by the cavity’s phase, and is independent of the qubit’s initial state. Consequently, the qubit is purified even if it was not initially in a pure state. Besides, an entropy decrease could also be induced in the qubit by simply coupling it to another system of low entropy, e.g. another qubit initially in a pure state. In contrast here, the cavity can have initially an entropy comparable to, or even higher than that of the qubit, while still enabling perfect purification, provided its coherent amplitude is large enough. This is possible owing to the infinite Hilbert space of the cavity, able to perfectly store the result of the measurement, and conserve it while performing a unitary feedback on the qubit. Another pre-requisite is the non-thermal nature of the initial cavity state, which makes it suitable to provide work [30]. Finally, notice that a large Hilbert space is also required for the ideal work source behavior observed at short times: more precisely, the work source Hilbert space needs to support quantum states of average energy differing by the qubit transition energy $\hbar\omega_{0}$, while also having an overlap arbitrarily close to $1$. This combination of properties, specific to infinite-dimensional Hilbert spaces, is required to generate coherent, purity preserving, Rabi oscillations of the qubit, despite the dynamical constraint of conserving the cavity and qubit total energy [48, 49].
IX Conclusions
@@
We have analyzed the dynamics of a qubit coupled to a cavity prepared in a displaced coherent state of very large displacement amplitude, a state commonly considered as guaranteeing a classical drive behavior of the cavity. Owing to our framework for the thermodynamics of autonomous machines, we have shown that at short enough times, the cavity indeed behaves as an ideal work source inducing a time-dependent drive on the qubit. However, over a time-scale set by the qubit-cavity interaction strength, the cavity performs a measurement of the qubit state in a basis set by the initial cavity displacement phase, responsible for the collapse of the Rabi oscillations. We have shown that this measurement is followed by an autonomous feedback, a state-dependent drive able to purify the qubit. From the analysis of thermodynamic quantities, we relate this purification to the cavity behaving as an autonomous Maxwell demon, spending work to build mutual information with the qubit, and then consuming this mutual information to decrease the qubit’s entropy. Those results demonstrate the potential of consistent thermodynamic frameworks suitable for autonomous systems, like the one of Ref. [30], to shed new light on complex quantum dynamics. The parameters needed to implement our scheme are within state-of-the-art, e.g. in superconducting qubit-cavity systems, where strong couplings with respect to damping rates can be achieved. Natural extensions of this work include optimizing the initial cavity state to improve the purification process, e.g. by attaining purer states at a fixed feedback time, and testing the robustness of the purified qubit state to cavity preparation and residual qubit-cavity detunings.
Acknowledgements
@@
C.E. acknowledges funding from French National Research Agency (ANR) under project ANR-22-CPJ1-0029-01. We thank Alexia Auffèves for fruitful discussions.
Appendix A Physical Model
@@
$\hat{H}$ $=\hat{H}_{\text{C}}+\hat{H}_{\text{Q}}+\hat{V}_{\text{QC}}=\hbar\omega_{0}a^{\dagger}a+\hbar\omega_{0}\sigma_{+}\sigma_{-}+i\hbar g\left(a\sigma_{+}-a^{\dagger}\sigma_{-}\right)$ (27)
@@
Due to the resonance condition, the evolution of the total system in the interaction picture is governed by Hamiltonian $\hat{H}_{I}(t)=\hat{V}_{\text{QC}}$. We deduce the interaction picture evolution operator $\hat{U}(t)=e^{-\frac{it}{\hbar}\hat{V}_{\text{QC}}}$, which can expanded as [38]:
$\hat{U}\left(t\right)=\cos\left(gt\sqrt{a^{\dagger}a+1}\right)|e\rangle\langle e|+\cos\left(gt\sqrt{a^{\dagger}a}\right)|g\rangle\langle g|+\frac{\sin\left({gt\sqrt{a^{\dagger}a+1}}\right)}{\sqrt{a^{\dagger}a+1}}a\sigma_{+}-\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}\sigma_{-}$ (28)
Appendix B Quasi-ideal work source behavior of the cavity
@@
$\hat{\rho}\left(0\right)=|e\rangle\langle e|\otimes D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right),\hskip 5.0pt\alpha_{0}=\sqrt{n_{0}}e^{i\phi_{0}}\;.$ (29)
@@
$\hat{\rho}\left(t\right)$ $=|e\rangle\langle e|\otimes\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)$ (30)
- $|e\rangle\langle g|\otimes\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}$
$-|g\rangle\langle e|\otimes\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)$
$+|g\rangle\langle g|\otimes\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}\;.$
@@
$\hat{\rho}_{\text{Q}}\left(t\right)=\begin{bmatrix}P_{e}\left(t\right)&C_{eg}\left(t\right)\\ C_{eg}\left(t\right)&1-P_{e}\left(t\right)\end{bmatrix}$ (31)
@@
$P_{e}\left(t\right)$ $=\sum_{n}\cos^{2}\left({gt\sqrt{n+1}}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ (32)
$C_{eg}\left(t\right)$ $=-\sum_{n}\cos\left({gt\sqrt{n+2}}\right)\sin\left({gt\sqrt{n+1}}\right)\langle n+1|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle\;.$
For $n_{0}\gg 1$ the terms $\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ and $\langle n+1|D\left(\alpha\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha\right)|n\rangle$ as a function of $n$ are close to Gaussian distributions peaked around $n_{0}$ with variance much smaller than the mean. This allows us to expand the trigonometric functions in $P_{e}\left(t\right)$ and $C_{eg}\left(t\right)$ around $n_{0}$. In the following, we parameterize the population and coherence of qubit in terms of rotation angle $\theta=2gt\sqrt{n_{0}}$ in the Bloch sphere.
@@
$P_{e}\left(t\right)$ $=\sum_{n}\cos^{2}\left({\frac{\theta}{2}\sqrt{1+\frac{n-n_{0}+1}{n_{0}}}}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$ (33)
@@
$\times\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)|n\rangle$
@@
$\delta P_{e}\left(\theta,\bar{n}\right)=-\frac{\theta}{16}\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)\;.$ (34)
@@
$C_{eg}\left(t\right)$ $=-e^{i\phi_{0}}[\frac{1}{2}\sin{\theta}+\frac{-2\theta+\left(3+2\bar{n}\right)\theta\cos\theta-\left(1+2\bar{n}\right)\left(1+\theta^{2}\right)\sin\theta}{16n_{0}}+$ (35)
@@
$=-e^{i\phi_{0}}[\frac{1}{2}\sin{\theta}+\frac{\delta C_{eg}\left(\theta,\bar{n}\right)}{n_{0}}+O\left(\frac{1}{n_{0}^{2}}\right)]$
@@
$\delta C_{eg}\left(\theta,\bar{n}\right)=\frac{-2\theta+\left(3+2\bar{n}\right)\theta\cos\theta-\left(1+2\bar{n}\right)\left(1+\theta^{2}\right)\sin\theta}{16}$ (36)
The expansions of $P_{e}(\theta)$ and $C_{eg}(\theta)$ can be used to find an expansion for the von Neumann entropy of the qubit up to first order in $n_{0}^{-1}$ in a similar way as
$S_{\text{Q}}(t)$ $=\epsilon_{\text{Q}}(\theta,\bar{n})\left(-\frac{\log\left(\epsilon_{\text{Q}}(\theta,\bar{n})/e\right)}{n_{0}}+\frac{\log n_{0}}{n_{0}}\right),$ (37)
where
$\epsilon_{\text{Q}}(\theta,\bar{n})$ $=\frac{(1+2\bar{n})(1+2\theta^{2}-\cos 2\theta)+4\theta\sin\theta}{32}.$ (38)
The term $\log n_{0}/{n_{0}}$ satisfies ${1}/{n_{0}}\leq\log n_{0}/{n_{0}}\leq 1/\sqrt{n_{0}}$, for $n_{0}\geq e$.